OpenStudy (anonymous):

A woman borrowed $10,000 at 12% interest compounded annually. If the loan was paid off in one lump sum at the end of 6 years, how much did she pay? is it the answer 17200 ?

OpenStudy (anonymous):

850000....

OpenStudy (anonymous):

10000 x (1.12)^6

OpenStudy (anonymous):

A = p(1+r)^t A = 10,000 ( 1.12^6) 10000 * 1.97 = $ 19,738.23

OpenStudy (anonymous):

what formula are you using? to solve this? i think there is Y=Ak^x or y=AB^x same formula just different variable and PMT=PV*(((r/n)/(1-(1+(r/n))^(-nt))

OpenStudy (anonymous):

ishan_meswani seems to bee the right answer

OpenStudy (anonymous):

i have no idea where jasb is getting his answers from i got the same as elecengineer and im sure its correct

OpenStudy (anonymous):

ishan_meswani I know this stuff it's fresh in my memory. since it's so easy to remember. And yes i know you are correct.

OpenStudy (anonymous):

i thought it was solve like this : interest in 6 years = 6x0.12 x 10000 = 7200 so the total gonna be interest + 10000 = 17200 ?

OpenStudy (anonymous):

it can't be that way. Since interest grows. and each year the total is different we could interpret it like this. 10000*1.12=11200 interest from year one. then the second year's interest would be multiplied by 11200*1.12=12544 same as 10000*1.12*1.12=12544

OpenStudy (anonymous):

i got it now , really appreciate

OpenStudy (anonymous):

you are welcome